OpenStudy (anonymous):

A team of 3 students is to be formed from a class of 25 students. How many ways are there to form this team if there is 1 team leader, 1 scribe and 1 timekeeper in the team and each person can only perform one role at a time?

OpenStudy (lgbasallote):

you use permutations here \[nPr = \frac{n!}{(n-r)!}\] therefore \[25P3 = \frac{25!}{(25-3)!}\] \[\implies \frac{25!}{22!}\]

OpenStudy (lgbasallote):

do you get it @ravin1991

OpenStudy (anonymous):

so my final answer i leave it it as 25!/22!

OpenStudy (lgbasallote):

no..you need to solve that

OpenStudy (anonymous):

hmm ok!

OpenStudy (anonymous):

thanks!